# Learning to Inference Adaptively for Multimodal Large Language Models

### v2

@@ Appendix A Further Implementation Details
Probabilistic execution plan sampling. Recall that in our probabilistic model, we define the distribution $p\left(\mathbf{s}|\{\mathbf{z}^{v|q}\},l,\mathbf{\phi}\right)$ via a sampling process. Given the input tokens and a latency budget $l$, the output of the lightweight scheduler is a logits vector corresponding to the $K$ available switches: $\pi_{1},\pi_{2},\dots\pi_{K}\in\mathbb{R}$, where $\pi_{i}$ represents the relative likelihood of selecting the $i^{\text{th}}$ switch. The latency budget $l$ allows us to define $k$, the maximum number of switches allowed to activate. Then, a sampled execution plan from $p\left(\mathbf{s}|\{\mathbf{z}^{v|q}\},l,\mathbf{\phi}\right)$ can be uniquely defined by a subset of $k$ distinct elements from $\{1,2,\dots,K\}$, corresponding to its activated switches. We sample the execution plan by randomly picking $k$ switches one by one, without replacement, following the logits $\{\pi_{i}\}_{i=1}^{K}$. The complete sampling procedure is summarized in Algorithm 1, where Cat$\left(\Omega,\{\pi_{i}:i\in\Omega\}\right)$ denotes the categorical distribution of selecting an element from $\Omega$ with probabilities parameterized by $\{\eta_{i}:i\in\Omega\}$ = Softmax($(\{\pi_{i}:i\in\Omega\})$. The process ensures that the sampled execution plan adheres to the input budget while maximizing the utilization.
@@
       Sample $\omega\sim$ Cat$\left(\Omega,\{\pi_{i}:i\in\Omega\}\right)$
@@
Differentiable sampling with Gumbel-Softmax. Our designed scheduler is difficult to train as it involves a non-differentiable discrete sampling process, which prevents gradients from backpropagate to the scheduler during training. A common workaround involves using a score function estimator [16, 59]; however, this method often suffers from high variance and slow convergence. Instead, we employ Gumbel-Softmax [22], a reparameterization trick for sampling from categorical distribution. In our implementation, the Gumbel-Softmax approximates $\omega\sim$ Cat$\left(\Omega,\{\pi_{i}:i\in\Omega\}\right)$ with a continuous random vector $\tilde{\omega}$:
$\tilde{\omega}=\text{Softmax}\left([g_{i}+\log\eta_{i}]\right)_{i\in\Omega},$ (8)
where each $g_{i}$ is i.i.d. sample drawn from Gumbel$(0,1)$; and $\eta_{i}=\text{Softmax}(\{\pi_{j}:j\in\Omega\})[i]$ is the probability of activating the $i^{\text{th}}$ switch, computed by the scheduler. Note that $\tilde{\omega}$ is continuous and has a well-defined gradient. To maintain a hard execution plan, we take the one-hot encoding of $\tilde{\omega}$ and apply the straight-through estimator (see [22] for more details).
@@ Appendix B Detailed Results
Full results on LLaVA 1.5. We report the full set of results on LLaVA 1.5, LLaVA-PruMerge, LLaVA-PruMerge+ and FastV in Tab. 5, as a complement to Tab. 1. All experiments follow the same setting as described in Sec. 4.1. These results confirm that our AdaLLaVA framework successfully adapts to LLaVA 1.5 across different backbone sizes, and can be further combined with recent token selection methods (PruMerge, PruMerge+ and FastV) to further enhance efficiency. We maintain comparable performance while improving efficiency across multiple benchmarks. Additionally, our analysis reveals how performance varies under different latency constraints, demonstrating our framework’s ability to trade between accuracy and latency.
Broader benchmarks. We extend our AdaLLaVA-L framework on broader benchmarks reported in [36], namely GQA [21], SEED-Bench [29], MM-Vet [62], LLaVa-WILD [35], and VizWiz [17] (see Tab. 4). The model shows comparable performance and adaptive ability under different latency budget. The results demonstrate the strong generalization of AdaLLaVA to a wide range of benchmarks.
Model GQA SEED-Bench MM-Vet LLaVa-WILD VizWiz
@@
Table 4: Results on broader benchmarks.
@@
Generalization across MLLMs. We demonstrate that AdaLLaVA can generalize to other MLLMs beyond LLaVA. We consider Mipha-3B [67], a lightweight MLLM built on Phi-2.7B [23]. Specifically, we apply AdaLLaVA-L on Mipha-3B, following its training strategy [67], and report the results on a comprehensive MLLM benchmark (MME), shown in Fig. 6. We see that AdaLLaVA-L maintains comparable performance under full computational budgets. With reduced compute budgets, AdaLLaVA-L shows minimal performance degradation: an average accuracy drop of only 3.4% at 85% budget and 6.1% at 60% budget. These results have similar trend to those with LLaVA-1.5 in Fig. 1.
Figure 6: Relative performance of applying AdaLLaVA-L to Mipha-3B under various latency budget. The center of the radar corresponds to 60% performance of the base Mipha-3B.
@@
Table 5: Results of MLLMs on six benchmarks. Our AdaLLaVA  can be applied to LLaVA 1.5 with different size of LLM with different design of switches. Percentage (%): The input latency requirement. AdaLLaVA-L: switches on selecting different transformer blocks. AdaLLaVA-H: switches on select different attention heads and MLP activations. VQAv2: VQAv2 set. SQAI: ScienceQA set. VQAT: TextVQA set. Prumerge: LLaVA 1.5 with PruMerge.
@@ Appendix C Additional Ablation Studies
Number & granularity of switches. We here conduct ablation studies to examine how the number and granularity of switches affect performance. Fig. 7 (Left) compares switches for the last 16 layers (used in Sec. 4) versus 24 layers in AdaLLaVA-L. While 24 switches enable finer FLOPs control, they significantly reduce model performance. The 16-switch configuration provides better accuracy while maintaining efficient adaptability. Fig. 7 (Right) evaluates attention sampling group sizes in AdaLLaVA-H, focusing on operations within the last 16 layers. While both 4-head and 8-head (used in Sec. 4) configurations show comparable performance-latency tradeoffs, the 4-head version enables more granular latency control.
Figure 7: Ablation studies on switch design choices.
@@
As shown in Fig. 7, from FLOPs ranging from 5T to 8T, AdaLLaVA-H-8-head shows slightly better performance overall, reaching approximately 76% on VQA v2 Accuracy compared to AdaLLaVA-L-16-switch which peaks around 75%. Moreover, AdaLLaVA-H demonstrates finer-grained control over the accuracy-latency trade-off. This is evident from the smoother curve of AdaLLaVa-H, which can be attributed to its head/neuron-level switches providing more granular control over computational resources compared to the layer-level switches. This flexibility allows AdaLLaVa-H to accommodate a wider range of latency budgets.
Appendix D Additional Results on Latency and Content Adaptivity
@@
Model Response under different latency. Here we show additional results on model response given same image-text input under different latency budget, similar to Fig. 1. As shown in Tab. 6, given an image-query pair and latency constraint, AdaLLaVA  learns to generate appropriate responses while adapting to varying computational budgets.
@@
Table 6: AdaLLaVA-L on LLaVA 1.5-7b model, generating appropriate responses while adapting to varying computational budgets.
@@
Fig. 8 further demonstrate the model’s content-aware behavior. For in the father-child scene, attention spreads across the entire street view for scene description but concentrates on the middle when asking about their activity. For the Happy Plaza image, attention focuses on the storefront sign when asking about the location name, but shifts to the promotional signage area when querying about special offers. Similarly, in the restaurant scene, attention distributes across the interior elements (tables, counter, chairs) when identifying the location type, but concentrates specifically on the woman’s clothing when asked about her attire. This consistently shows how the model adjusts its attention based on query requirements.
Figure 8: The key-query attention scores between latency token and visual tokens. The latency input is 1.0 in these examples.
Visualization for latency token across layers. We plot the evolution of the latency token from layers 12 to 16 using the same example in Fig. 4 of the main paper (see figure below). As seen in Fig. 9, the latency token progressively gathers key information from the input visual tokens for scheduling.
Figure 9: Evolution of latency token across layers in AdaLLaVA-L on 7b model.

### v3

@@ Appendix A Further Implementation Details
[…]
@@
[…]
@@
[…]
$\tilde{\omega}=\text{Softmax}\left([g_{i}+\log\eta_{i}]\right)_{i\in\Omega},$ (A)
where each $g_{i}$ is i.i.d. sample drawn from Gumbel(0,1)(0,1); and $\eta_{i}=\text{Softmax}(\{\pi_{j}:j\in\Omega\})[i]$ is the probability of activating the $i^{\text{th}}$ switch, computed by the scheduler. Note that $\tilde{\omega}$ is continuous and has a well-defined gradient. To maintain a hard execution plan, we take the one-hot encoding of $\tilde{\omega}$ and apply the straight-through estimator (see [22] for more details).
Training details. Training details were discussed in Sec. 4 of the main paper. Here we show training curve of AdaLLaVA-L-7B with LLaVA 1.5 in Fig. A.
Figure A: Training logs
@@ Appendix B Detailed Results
Full results on LLaVA 1.5. We report the full set of results on LLaVA 1.5, LLaVA-PruMerge, LLaVA-PruMerge+ and FastV in Tab. B, as a complement to Tab. 1. All experiments follow the same setting as described in Sec. 4.1. These results confirm that our AdaLLaVA framework successfully adapts to LLaVA 1.5 across different backbone sizes, and can be further combined with recent token selection methods (PruMerge, PruMerge+ and FastV) to further enhance efficiency. We maintain comparable performance while improving efficiency across multiple benchmarks. Additionally, our analysis reveals how performance varies under different latency constraints, demonstrating our framework’s ability to trade between accuracy and latency.
Broader benchmarks. We extend our AdaLLaVA-L framework on broader benchmarks reported in [36], namely GQA [21], SEED-Bench [29], MM-Vet [64], LLaVa-WILD [35], and VizWiz [17] (see Tab. A). The model shows comparable performance and adaptive ability under different latency budget. The results demonstrate the strong generalization of AdaLLaVA to a wide range of benchmarks.
Model GQA SEED-Bench $M$-Vet LLaVa-WILD VizWiz
@@
Table A: Results on broader benchmarks.
@@
Generalization across MLLMs. We demonstrate that AdaLLaVA can generalize to other MLLMs beyond LLaVA. We consider Mipha-3B [70], a lightweight MLLM built on Phi-2.7B [23]. Specifically, we apply AdaLLaVA-L on Mipha-3B, following its training strategy [70], and report the results on a comprehensive MLLM benchmark (MME), shown in Fig. B. We see that AdaLLaVA-L maintains comparable performance under full computational budgets. With reduced compute budgets, AdaLLaVA-L shows minimal performance degradation: an average accuracy drop of only 3.4% at 85% budget and 6.1% at 60% budget. These results have similar trend to those with LLaVA-1.5 in Fig. 1.
Figure B: Relative performance of applying AdaLLaVA-L to Mipha-3B under various latency budget. The center of the radar corresponds to 60% performance of the base Mipha-3B.
@@
Table B: Results of MLLMs on six benchmarks. Our AdaLLaVA  can be applied to LLaVA 1.5 with different size of LLM with different design of switches. Percentage (%): The input latency requirement. AdaLLaVA-L: switches on selecting different transformer blocks. AdaLLaVA-H: switches on select different attention heads and MLP activations. VQAv2: VQAv2 set. SQAI: ScienceQA set. VQAT: TextVQA set. Prumerge: LLaVA 1.5 with PruMerge.
@@ Appendix C Additional Ablation Studies
Number & granularity of switches. We here conduct ablation studies to examine how the number and granularity of switches affect performance. Fig. C (Left) compares switches for the last 16 layers (used in Sec. 4) versus 24 layers in AdaLLaVA-L. While 24 switches enable finer FLOPs control, they significantly reduce model performance. The 16-switch configuration provides better accuracy while maintaining efficient adaptability. Fig. C (Right) evaluates attention sampling group sizes in AdaLLaVA-H, focusing on operations within the last 16 layers. While both 4-head and 8-head (used in Sec. 4) configurations show comparable performance-latency tradeoffs, the 4-head version enables more granular latency control.
Figure C: Ablation studies on switch design choices.
@@
As shown in Fig. C, from FLOPs ranging from 5T to 8T, AdaLLaVA-H-8-head shows slightly better performance overall, reaching approximately 76% on VQA v2 Accuracy compared to AdaLLaVA-L-16-switch which peaks around 75%. Moreover, AdaLLaVA-H demonstrates finer-grained control over the accuracy-latency trade-off. This is evident from the smoother curve of AdaLLaVA-H, which can be attributed to its head/neuron-level switches providing more granular control over computational resources compared to the layer-level switches. This flexibility allows AdaLLaVA-H to accommodate a wider range of latency budgets.
Comparison with naive sampling strategies. We compare the performance of our AdaLLaVA-L versus random uniform sampling, where we disable the scheduler during training, showing in Fig. D. Both methods are built on Mipha-3B and fine-tuned using the same procedure. Random sampling is worse than AdaLLaVA and has high variance in results (shaded area).
Figure D: Comparison to random sampling
Figure E: Visualization of execution plan
Appendix D Additional Results on Adaptivity
@@
Model Response under different latency. Here we show additional results on model response given same image-text input under different latency budget, similar to Fig. 1. As shown in Tab. C, given an image-query pair and latency constraint, AdaLLaVA  learns to generate appropriate responses while adapting to varying computational budgets.
Visualization of execution plans with different latency. We report execution plans of the same input with varying budgets in Fig. E. As budget decreases, the scheduler prioritizes keeping the last 3 layers over others. We have included outputs for the same input with varying budgets in Tab. C.
@@
Table C: AdaLLaVA-L on LLaVA 1.5-7b model, generating appropriate responses while adapting to varying computational budgets.
@@
Fig. F further demonstrate the model’s content-aware behavior. In the father-child scene image, attention spans the entire street for scene description but focuses centrally for query asking for activity. For Happy Plaza image, attention targets the storefront sign for location queries but shifts to promotional areas for query about special offers. In the restaurant scene, attention distributes across interior elements when identifying location type but concentrates on the woman’s clothing for attire questions. This consistently shows model adjusts its attention based on the query.
Figure F: The key-query attention scores between latency token and visual tokens. The latency input is 1.0 in these examples.
Visualization for latency token across layers. We plot the evolution of the latency token from layers 12 to 16 using the same example in Fig. 4 of the main paper (see figure below). As seen in Fig. G, the latency token progressively gathers key information from the input visual tokens for scheduling.
Figure G: Evolution of latency token across layers in AdaLLaVA-L on 7b model.
Appendix E Further Discussion
FLOPs, latency, and cross-device portability. In our work, compute budgets are expressed as percentages of a base model’s FLOPs. Percentages can be translated into absolute FLOP targets, given base model’s architecture and input size, from which total FLOPs can be estimated. We choose FLOPs over runtime latency, as it abstracts away hardware / software specific variations . We acknowledge that the relationship between FLOPs and latency vary across devices, depending on model architecture, hardware and software. We leave cross-device portability to future work.